OpenStudy (anonymous):

A certain product is packed in cubic boxes that measure 2 feet on a side. If a rectangular shipping container is 6 feet long, 4 feet wide, and 8 feet high, how many of these cubic boxes will be able to fit inside it? 1. 24 2. 16 3. 8 4. 4 5. 2

OpenStudy (austinl):

\(\text{Volume}_{Big~Box}=l\times w\times h\) \(\text{Volume}_{Little~Box}=l\times w\times h\) \(Total~Number~of~Boxes = \dfrac{\text{Volume}_{Big~Box}}{\text{Volume}_{Little~Box}}\)

OpenStudy (anonymous):

so 2

OpenStudy (austinl):

No, \(\large{\dfrac{6\times4 \times 8}{2\times 2\times2}}\)

OpenStudy (anonymous):

24